# Multi-Frequency Progressive Refinement for Learned Inverse Scattering

### v1

@@ 5.2 Alternative Multi-Frequency Methods
We wish to compare our method, MFISNet-Refinement, with other methods of learning an inverse to the multi-frequency forward map. We design two new methods of extending FBP-inspired single-frequency architectures to the multi-frequency setting. We use the FYNet architecture (Fan and Ying, 2022) to instantiate all three of our MFISNet methods, which allows us to focus on the effects caused by different methods of combining multi-frequency data. We show the architectures in Figure 4. We also compare our method with the Wide-Band Butterfly Network (Li et al., 2022, 2021a).
@@ 5.3 Stabilizing Reconstruction by Adding Frequencies
For the $N_{k}=1$ dataset, we train an FYNet model, and for the other datasets, we train our three models: MFISNet-Fused, MFISNet-Parallel, and MFISNet-Refinement. For each model, we train for a fixed number of epochs and choose the model weights at the epoch at which a validation set of size $n/10$ is minimized. We also use the validation set to search over various hyperparameters, such as the size of 1D and 2D convolutional kernels, the number of channels in the convolutional layers, and optimization hyperparameters, such as step size and weight decay. For the Wide-Band Butterfly Network, we search over the rank of the butterfly factorization, as well as optimization hyperparameters, such as initial learning rate, batch size, and learning rate decay (Appendix B.)
We present the quantitative results of this experiment in Table 1 and Figure 6 (see also Figures 5 and 8 for a qualitative comparison on representative test samples). The relatively poor performance of FYNet confirms our belief that we are in a challenging nonlinear problem regime. As more frequencies are added, the multi-frequency methods improve. The performances of MFISNet-Fused and MFISNet-Parallel are comparable, indicating that the distinction between learning adjoint operators separately or jointly does not have a large effect in this setting. All of tested methods are uniformly outperformed by our method, MFISNet-Refinement, at all values of $N_{k}$.
@@
[…]
[width=0.75]svg-inkscape/preds_compared_along_n_omega_redo_0_svg-tex.pdf_tex
Figure 5: Sample predictions from models trained on different datasets. Predictions and errors on one held-out test sample are shown. The first row shows the ground-truth scattering potential, as well as the predictions and errors from the single-frequency baseline method, FYNet. The second and third rows show predictions and errors, respectively, from MFISNet models trained on $N_{k}=3$ frequencies. The fourth and fifth rows show predictions and errors, respectively, from MFISNet models trained on $N_{k}=5$ frequencies. See Figure 8 in the appendix for additional samples and outputs from Wide-Band Butterfly Network.
@@
1 $[32\pi]$ $10,000$ FYNet $0.261\pm 0.036$
2 $[16\pi,32\pi]$ $5,000$ MFISNet-Fused $0.177\pm 0.033$
MFISNet-Parallel $0.168\pm 0.029$
MFISNet-Refinement (Ours) $\mathbf{0.154\pm 0.034}$
3 $[8\pi,16\pi,32\pi]$ $33,3$ 333 Wide-Band Butterfly Network $0.160\pm 0.037$
MFISNet-Fused $0.130\pm 0.025$
MFISNet-Parallel $0.114\pm 0.021$
MFISNet-Refinement (Ours) $\mathbf{0.098\pm 0.020}$
4 $[4\pi,8\pi,16\pi,32\pi]$ $2,500$ MFISNet-Fused $0.105\pm 0.021$
MFISNet-Parallel $0.110\pm 0.021$
MFISNet-Refinement (Ours) $\mathbf{0.086\pm 0.017}$
@@
MFISNet-Parallel $0.108\pm 0.021$
MFISNet-Refinement (Ours) $\mathbf{0.084\pm 0.018}$
@@
$32\pi$$16\pi,32\pi$$8\pi,16\pi,32\pi$$4\pi,8\pi,16\pi,32\pi$$2\pi,4\pi,8\pi,16\pi,32\pi$00.050.10.150.20.250.3Input frequenciesRelative $\ell_{2}$ errorFYNetWide-Band Butterfly NetMFISNet-FusedMFISNet-ParallelMFISNet-Refinement (Ours)
Figure 6: Illustration of the results of Table 1. The advantage of the Residual FYNet model grows as more low-frequency data is used, even as the total volume of training data is held constant regardless of the number of frequencies.
@@ 5.4 Measurement Noise
Under this noise model, the expected noise-to-signal ratio is $\mathbb{E}_{Z\sim\mathcal{N}(0,I)}[\|\tilde{d}_{k}-d_{k}\|/\|d_{k}\|]\approx\delta$. We do not alter the scattering potentials $q_{i}$ in any way. We train and test all models using noisy inputs, and report the results of this experiment in Table 2. For the Wide-Band Butterfly Network, which takes $d_{k}$ inputs in the original $(r,s)$ coordinates, we add noise to the input in its original coordinates and divide by $\sqrt{2N_{r}N_{s}}$ instead. This results in an equivalent noise profile thanks to the normalization that is grid-invariant. We note that the models lose between $1\%$ $2\%$ accuracy on test set, suggesting the methods are relatively robust to the presence of measurement noise. Again, our method, MFISNet-Refinement, outperforms all other methods at all values $N_{k}$. We present the results of this experiment in Table 2 and visualize the results in Figure 7.
@@
1 $[32\pi]$ $10,000$ FYNet $0.251\pm 0.037$
2 $[16\pi,32\pi]$ $5,000$ MFISNet-Fused $0.170\pm 0.033$
MFISNet-Parallel $0.169\pm 0.031$
MFISNet-Refinement (Ours) $\bm{0.155\pm 0.034}$
3 $[8\pi,16\pi,32\pi]$ $33,3$ 333 Wide-Band Butterfly Network $0.159\pm 0.037$
MFISNet-Fused $0.135\pm 0.025$
MFISNet-Parallel $0.116\pm 0.024$
MFISNet-Refinement (Ours) $\bm{0.100\pm 0.021}$
4 $[4\pi,8\pi,16\pi,32\pi]$ $2,500$ MFISNet-Fused $0.107\pm 0.020$
MFISNet-Parallel $0.105\pm 0.020$
MFISNet-Refinement (Ours) $\bm{0.087\pm 0.017}$
5 $[2\pi,4\pi,8\pi,16\pi,32\pi]$ $2,000$ MFISNet-Fused $0.118\pm 0.024$
MFISNet-Parallel $0.110\pm 0.021$
MFISNet-Refinement (Ours) $\bm{0.095\pm 0.018}$
@@
$32\pi$$16\pi,32\pi$$8\pi,16\pi,32\pi$$4\pi,8\pi,16\pi,32\pi$$2\pi,4\pi,8\pi,16\pi,32\pi$00.050.10.150.20.250.3Input frequenciesRelative $\ell_{2}$ errorFYNetWide-Band Butterfly NetMFISNet-FusedMFISNet-ParallelMFISNet-Refinement (Ours)
@@ 5.5 Investigating the Training Method
We hypothesize that our method is successful because it effectively breaks the reconstruction problem into a sequence of simpler frequency-dependent refinement steps. This emulates the sequential, frequency-dependent structure of the recursive linearization method. A natural question is whether the success of our method is primarily driven by emulating sequential refinement steps or can be attributed to the residual architecture alone. To test this question, we investigate two different changes to our training method which remove the progressive refinement structure. We describe each adjustment below.
@@
Instead of using the intermediate loss terms $\|\hat{q}_{k_{t}}-\mathsf{LPF}_{2k_{t}}q\|^{2}$, designed to promote specific network blocks learning different parts of the reconstruction, we train the network by only optimizing the final loss term $\|\hat{q}_{k_{N_{k}}}-q\|^{2}$. This is the standard training loss used in most other works, including Khoo and Ying (2019); Fan and Ying (2022); Li et al. (2022).
Training Method ${[k_{1},k_{2},...]}$ Relative L2 Error
No Progressive Refinement $[2\pi,4\pi,8\pi,16\pi,32\pi]$ $0.095\pm 0.018$
No Homotopy through Frequency $[2\pi,4\pi,8\pi,16\pi,32\pi]$ $0.090\pm 0.017$
Our Method $[2\pi,4\pi,8\pi,16\pi,32\pi]$ $\mathbf{0.084\pm 0.018}$
Table 3: Our training method, described in Algorithm 2, produces better results than the other methods, which were designed to remove parts of the recursive linearization structure. We describe these alternate training methods in Section 5.5.
@@ 6 Conclusion
This work also leaves open important questions about machine learning in different multi-frequency data settings. We leave the investigation of these important problems to future work. The first setting is seismic imaging, where sources and receivers are located on one side of the scattering potential, resulting in limited-aperture measurements. Another setting to consider is full-aperture measurements, with a small fixed or frequency-dependent number of source and receiver directions, as $N_{s}$ and $N_{r}$ drive real-world costs when implementing an imaging system. Finally, we suggest a distribution of scattering potentials with an unknown, smoothly varying background occluded by strongly scattering shapes. We note that an important open problem is to learn to segment the reconstruction into disjoint regions, containing only background or only the strong scatterers.

### v2

@@ 5.2 Alternative Multi-Frequency Methods
We wish to compare our method, MFISNet-Refinement, with other methods of learning an inverse to the multi-frequency forward map. We design two new methods of extending FBP-inspired single-frequency architectures to the multi-frequency setting. We use the FYNet architecture (Fan and Ying, 2022) to instantiate all three of our MFISNet methods, which allows us to focus on the effects caused by different methods of combining multi-frequency data. We show the architectures in Figure 5. We also compare our method with the Wide-Band Butterfly Network (Li et al., 2022, 2021a). For broader context, we refer to Li et al. (2022) for comparisons between the Wide-Band Butterfly Network and classical multi-frequency methods that do not involve any machine learning such as Full Waveform Inversion (FWI) and a Least-Squares (LS) scheme. The authors show that the Wide-Band Butterfly Network achieves better accuracy than either FWI or LS with much faster inference times with minimal frequency or hyperparameter tuning.
[…]
@@ 5.3 Stabilizing Reconstruction by Adding Frequencies
For the $N_{k}=1$ dataset, we train an FYNet model, and for the other datasets, we train our three models: MFISNet-Fused, MFISNet-Parallel, and MFISNet-Refinement. For each model, we train for a fixed number of epochs and choose the model weights at the epoch at which the error on a validation set of size $n/10$ is minimized. We also use the validation set to search over various hyperparameters, such as the size of 1D and 2D convolutional kernels, the number of channels in the convolutional layers, and optimization hyperparameters, such as step size and weight decay. For the Wide-Band Butterfly Network, we search over the rank of the butterfly factorization, as well as optimization hyperparameters, such as initial learning rate, batch size, and learning rate decay (Appendix B).
We present the results of this experiment in Table 1 and Figures 6 and 7. See also Appendix D for additional empirical results, which include training and testing runtimes, visualizations of predictions on more held-out test samples, and visualizations of predictions of MFISNet-Parallel and MFISNet-Fused. The relatively poor performance of FYNet confirms our belief that we are in a challenging nonlinear problem regime. As more frequencies are added, the multi-frequency methods improve. The performances of MFISNet-Fused and MFISNet-Parallel are comparable, indicating that the distinction between learning adjoint operators separately or jointly does not have a large effect in this setting. For $N_{k}\geq 3$, the tested methods are uniformly outperformed by our method, MFISNet-Refinement. As we keep increasing $N_{k}$, the performance of MFISNet-Refinement plateaus. We hypothesize that our method could be improved by choosing a different set of incident wave frequencies, possibly linearly-spaced in a smaller frequency band. The optimal choice of frequencies could also be learned from data in a reinforcement learning setting (Jiang et al., 2024).
@@
Figure 6: Sample predictions from models trained on different datasets. Predictions and errors on two held-out test samples are shown. The first row shows the ground-truth scattering potential; in this plot we show the wavelength corresponding to the maximum frequency $k=32\pi$. The remaining rows show predictions and errors for FYNet, Wide-Band Butterfly Network, MFISNet-Refinement $(N_{k}=3)$, and MFISNet-Refinement $(N_{k}=5)$. See Appendix D for additional samples and outputs from MFISNet-Parallel and MFISNet-Fused.
@@
1 $[32\pi]$ $10,000$ FYNet $0.159\pm 0.033$
2 $[16\pi,32\pi]$ $5,000$ MFISNet-Fused $0.158\pm 0.030$
MFISNet-Parallel $\mathbf{0.144\pm 0.029}$
MFISNet-Refinement (Ours) $0.152\pm 0.032$
3 $[8\pi,16\pi,32\pi]$ $33,3$ 333 Wide-Band Butterfly Network $0.156\pm 0.037$
MFISNet-Fused $0.121\pm 0.024$
MFISNet-Parallel $0.107\pm 0.021$
MFISNet-Refinement (Ours) $\mathbf{0.094\pm 0.018}$
4 $[4\pi,8\pi,16\pi,32\pi]$ $2,500$ MFISNet-Fused $0.103\pm 0.020$
MFISNet-Parallel $0.106\pm 0.021$
MFISNet-Refinement (Ours) $\mathbf{0.082\pm 0.019}$
@@
MFISNet-Parallel $0.109\pm 0.021$
MFISNet-Refinement (Ours) $\mathbf{0.087\pm 0.018}$
@@
$32\pi$$16\pi,32\pi$$8\pi,16\pi,32\pi$$4\pi,8\pi,16\pi,32\pi$$2\pi,4\pi,8\pi,16\pi,32\pi$00.050.10.150.20.25Input frequenciesRelative $\ell_{2}$ errorFYNetWide-Band Butterfly NetMFISNet-FusedMFISNet-ParallelMFISNet-Refinement (Ours)
Figure 7: Illustration of the results of Table 1. The advantage of the MFISNet-Refinement model grows as more low-frequency data is used, even as the total volume of training data is held constant regardless of the number of frequencies.
$N_{k}$ $n$ Method Name Training Time Testing Time
(seconds) (seconds)
1 $10,000$ FYNet $870.3$ $27.3$
3 $33,3$ 333 Wide-Band Butterfly Network $6,453.6$ $29.9$
MFISNet-Fused $326.2$ $27.8$
MFISNet-Parallel $256.1$ $27.8$
MFISNet-Refinement (Ours) $2,033.0$ $28.2$
5 $2,000$ MFISNet-Fused $306.3$ $27.6$
MFISNet-Parallel $182.2$ $27.8$
MFISNet-Refinement (Ours) $2,095.5$ $28.8$
Table 2: The time required to train and evaluate machine learning models. In different settings of $N_{k}$, we report the training time for each method. We also report the time required to make predictions on the entire set of $1,000$ held-out test samples, performed in ten batches of size $100$. As we increase $N_{k}$, the number of samples $n$ decreases, so MFISNet-Fused and MFISNet-Parallel see an increase in training speed. For MFISNet-Refinement, as we increase the number of frequencies, we also increase the number of training epochs. The two effects, smaller dataset and more epochs, approximately negate each other, so the training time for MFISNet-Refinement remains approximately constant for different $N_{k}$. The testing times for the FYNet and MFISNet models are all approximately equal because they are dominated by a final polar-to-Cartesian coordinate transformation step, which is performed on a single CPU core but can in principle be accelerated.
@@ 5.4 Measurement Noise
Under this noise model, the expected noise-to-signal ratio is $\mathbb{E}_{Z\sim\mathcal{N}(0,I)}[\|\tilde{d}_{k}-d_{k}\|/\|d_{k}\|]\approx\delta$. We do not alter the scattering potentials $q_{i}$ in any way. We train and test all models using noisy inputs, and report the results of this experiment in Table 3. For the Wide-Band Butterfly Network, which takes $d_{k}$ inputs in the original $(r,s)$ coordinates, we add noise to the input in its original coordinates and divide by $\sqrt{2N_{r}N_{s}}$ instead. This results in an equivalent noise profile thanks to the normalization that is grid-invariant. We note that the models lose between $1\%$ $2\%$ accuracy on test set, suggesting the methods are relatively robust to the presence of measurement noise. Again, our method, MFISNet-Refinement, outperforms all other methods at most values $N_{k}$. We present the results of this experiment in Table 3 and visualize the results in Figure 8.
@@
1 $[32\pi]$ $10,000$ FYNet $0.163\pm 0.031$
2 $[16\pi,32\pi]$ $5,000$ MFISNet-Fused $0.164\pm 0.035$
MFISNet-Parallel $\mathbf{0.148\pm 0.029}$
MFISNet-Refinement (Ours) $0.151\pm 0.031$
3 $[8\pi,16\pi,32\pi]$ $33,3$ 333 Wide-Band Butterfly Network $0.156\pm 0.037$
MFISNet-Fused $0.125\pm 0.025$
MFISNet-Parallel $0.110\pm 0.023$
MFISNet-Refinement (Ours) $\mathbf{0.097\pm 0.019}$
4 $[4\pi,8\pi,16\pi,32\pi]$ $2,500$ MFISNet-Fused $0.102\pm 0.021$
MFISNet-Parallel $0.108\pm 0.019$
MFISNet-Refinement (Ours) $\mathbf{0.086\pm 0.019}$
5 $[2\pi,4\pi,8\pi,16\pi,32\pi]$ $2,000$ MFISNet-Fused $0.112\pm 0.022$
MFISNet-Parallel $0.103\pm 0.020$
MFISNet-Refinement (Ours) $\mathbf{0.090\pm 0.017}$
@@
$32\pi$$16\pi,32\pi$$8\pi,16\pi,32\pi$$4\pi,8\pi,16\pi,32\pi$$2\pi,4\pi,8\pi,16\pi,32\pi$00.050.10.150.20.25Input frequenciesRelative $\ell_{2}$ errorFYNetWide-Band Butterfly NetMFISNet-FusedMFISNet-ParallelMFISNet-Refinement (Ours)
@@ 5.5 Investigating the Training Method
Our method unfolds the reconstruction problem into a sequence of simpler frequency-dependent refinement steps. This emulates the sequential, frequency-dependent structure of the recursive linearization method. Compared to existing methods, our method introduces both a new residual architecture and a new sequential training procedure, so a natural question would be how important each of these factors is to our method’s success. To test this question, we investigate two different changes to our training method which remove the progressive refinement structure. We describe each adjustment below. The results are presented in Table 4. In this experiment, we find the accuracy of our method is relatively robust to perturbations of the training method presented in Algorithm 2. This indicates the advantage of our approach is primarily driven by the residual architecture.
@@
Instead of using the intermediate loss terms $\|\hat{q}_{k_{t}}-\mathsf{LPF}_{2k_{t}}q\|_{2}^{2}$, designed to promote specific network blocks learning different parts of the reconstruction, we train the network by only optimizing the final loss term $\|\hat{q}_{k_{N_{k}}}-q\|_{2}^{2}$. This is the standard training loss used in most other works, including Khoo and Ying (2019); Fan and Ying (2022); Li et al. (2022).
Training Method ${[k_{1},k_{2},...]}$ $n$ Relative L2 Error
No Progressive Refinement $[16\pi,32\pi]$ $5,000$ $\mathbf{0.152\pm 0.030}$
No Homotopy through Frequency $\mathbf{0.152\pm 0.030}$
Algorithm 2 $\mathbf{0.152\pm 0.032}$
No Progressive Refinement $[8\pi,16\pi,32\pi]$ $33,3$ 333 $\mathbf{0.094\pm 0.019}$
No Homotopy through Frequency $0.097\pm 0.018$
Algorithm 2 $\mathbf{0.094\pm 0.018}$
No Progressive Refinement $[4\pi,8\pi,16\pi,32\pi]$ $2,500$ $0.091\pm 0.020$
No Homotopy through Frequency $0.086\pm 0.018$
Algorithm 2 $\mathbf{0.082\pm 0.019}$
No Progressive Refinement $[2\pi,4\pi,8\pi,16\pi,32\pi]$ $2,000$ $\mathbf{0.082\pm 0.018}$
No Homotopy through Frequency $0.087\pm 0.016$
Algorithm 2 $0.087\pm 0.018$
Table 4: Alternative training methods, designed to remove parts of the recursive linearization structure, produce models of similar accuracy to the training procedure discussed in Algorithm 2. This indicates the advantage of our MFISNet-Refinement method is driven primarily by its residual architecture. We describe the alternate training methods in Section 5.5. The optimal hyperparameters for these models are listed in Appendix B.
5.6 Investigating the Speed of Training Convergence
While Tables 1 and 3 show that MFISNet-Refinement is more accurate than other networks in most settings, Table 2 shows that training our network is slower than the baseline MFISNet-Parallel and MFISNet-Fused architectures. This is because our MFISNet-Refinement architecture is more complex than MFISNet-Parallel or MFISNet-Fused, and our training procedure (Algorithm 2) takes a block-wise approach which requires many training epochs. In this section, we conduct experiments to investigate whether we can reduce the time spent training MFISNet-Refinement. Because the results presented in Section 5.5 suggest that all three training methods considered, Algorithm 2, “No Homotopy through Frequency”, and “No Progressive Refinement”, produce models which are approximately similar in accuracy, we consider training with all of these methods.
We introduce another training method, designed to mimic Algorithm 2 but converge in fewer training epochs. This method, which we call “Algorithm 2 + Warm-Start Initialization”, uses the previous block’s parameters $\theta_{t-1}$ to initialize the weights of the next block $\theta_{t}$. This “warm start” is inspired by the fact that neighboring blocks in our MFISNet-Refinement architecture are designed to learn similar functions; each block is meant to increase the resolution of the estimated scattering potential by a small amount. Thus, the weights learned by the previous block may be a good starting point for the optimization of the next block. We give full pseudocode for this algorithm in Appendix C.
We set a early stopping criterion, which terminates training if the relative $\ell_{2}$ error has not decreased by at least $10^{-3}$ over the past 15 epochs. We use this early stopping condition to terminate the “No Homotopy through Frequency” and “No Progressive Refinement” training methods. Similarly, we use this early stopping criterion to terminate each block in Algorithm 2. In Table 5, we measure the number of epochs used and runtime of each training method. We observe that the warm-start initialization decreases the number of epochs and training time of Algorithm 2. For $N_{k}=3$ frequencies, this method converges faster than the other three methods. In both cases, using the warm start initialization incurs a slight decrease in accuracy while greatly increasing the speed of training convergence. Different early stopping criteria could achieve different tradeoffs between training speed and accuracy, and we leave the investigation of such criteria to future work.
The time required to complete an epoch of training for Algorithm 2 with and without warm-starting is much shorter than that of the “No Progressive Refinement” and “No Homotopy through Frequency” methods. This is because for the majority of epochs in Algorithm 2, only one block is updated at a time, making the backpropogation of gradients much faster. In the “No Progressive Refinement” and “No Homotopy through Frequency” methods, all trainable parameters are updated in each epoch.
Training Method ${[k_{1},k_{2},...]}$ $n$ Number of Training Time Relative $\ell_{2}$
Epochs (seconds) Error
No Progressive Refinement $[8\pi,16\pi,32\pi]$ $33,3$ 333 $90$ $1,200.6$ $0.096\pm 0.020$
No Homotopy through Frequency $55$ $920.3$ $0.105\pm 0.020$
Algorithm 2 $215$ $1,226.4$ $0.094\pm 0.019$
Algorithm 2 + Warm-Start Init. $195$ $802.8$ $0.098\pm 0.019$
No Progressive Refinement $[2\pi,4\pi,8\pi,16\pi,32\pi]$ $2,000$ $45$ $811.6$ $0.095\pm 0.019$
No Homotopy through Frequency $50$ $997.8$ $0.091\pm 0.017$
Algorithm 2 $260$ $1,417.5$ $0.087\pm 0.017$
Algorithm 2 + Warm-Start Init. $210$ $1041.9$ $0.105\pm 0.020$
Table 5: Using warm-starting in Algorithm 2 reduces the training runtime between $20$ and $35\%$. In this table, we report time and epochs required to train MFISNet-Refinement  models with different training methods, and the relative $\ell_{2}$ errors of each fully-trained model. The warm-start initialization method and the convergence criterion are described in Section 5.6.
@@ 6 Conclusion
This work also leaves open important questions about machine learning in different multi-frequency data settings. We leave the investigation of these important problems to future work. The first setting is seismic imaging, where sources and receivers are located on one side of the scattering potential, resulting in limited-aperture measurements. Another setting to consider is full-aperture measurements, with a small fixed or frequency-dependent number of source and receiver directions, as $N_{s}$ and $N_{r}$ drive real-world costs when implementing an imaging system. Real-world imaging systems often encounter noise that is not well-approximated by an additive zero-mean Gaussian noise model; characterizing the robustness of deep learning methods in more realistic noise settings is important future work. Finally, we suggest a distribution of scattering potentials with an unknown, smoothly varying background occluded by strongly scattering shapes. We note that an important open problem is to learn to segment the reconstruction into disjoint regions, containing only background or only the strong scatterers.
@@
Data Availability
Our publicly-available GitHub repository https://github.com/meliao/mfisnets contains the following:
  • Code for defining and training MFISNet-Refinement, MFISNet-Parallel, and MFISNet-Fused.
  • Code for generating the dataset used in our experiments.
Our publicly-available dataset of $10,000$ training samples, $1,000$ validation samples, and $1,000$ test samples can be downloaded from https://doi.org/10.5281/zenodo.14514353.